bata =4 is found in quadrant 3 and bata= ?
Answer by stanbon(75887)

beta = 4 is found in quadrant 3 and
beta= ?
----------
4-pi = 4-3.14159 = 0.8584
4 is in Quadrant 3
======================
Cheers,
Stan H.
=================